# Boaz Moerman

### v1

@@ Abstract.
We establish an asymptotic formula for the number of $\mathcal{M}$-points of bounded height on split toric varieties, for the height induced by any big and nef divisor class. This formula establishes new cases of the extension of Manin’s conjecture to $\mathcal{M}$-points, as introduced by the author. As a special case of our result, we strengthen the results obtained by Pieropan and Schindler on Campana points of bounded height on toric varieties. As another special case, we obtain an asymptotic for the number of weak Campana points of bounded height, which is novel even for projective space.
We illustrate our result by giving an asymptotic for the number of points on projective space of bounded height for which the product of coordinates is powerful.
@@ 1. Introduction
In this article, we obtain an asymptotic for the number of $\mathcal{M}$-points of bounded height on split toric varieties, in the spirit of Manin’s conjecture. Since its introduction in the 1980’s [FMT89], Manin’s conjecture has inspired many developments in arithmetic geometry and arithmetic statistics, and it has recieved various refinements [BaMa90, Pey95, BaTs98, LeSeKuSh22]. This conjecture predicts a precise asymptotic for the number of rational points of bounded height on rationally connected varieties.
A few years ago, Manin’s conjecture was extended by Pieropan, Smeets, Tanimoto and Várilly-Alvarado to a conjecture on Campana points of bounded height [PSTVA21]. This conjecture has attracted a lot of interest, and various cases of it have been proven [Val12, Xia21, BBKOPW23, PiSc23, PiSc24, CLTT24]. In a recent work [Moe25conjecture], the author further extended Manin’s conjecture to the broader setting of $\mathcal{M}$-points. This conjecture gives a general framework for various results on Campana points and Darmon points which fall outside of the scope of the previous conjectures [Str21, ShSt24, Ito25, Ara25]. In this article, we will prove the latter conjecture for $\mathcal{M}$-points of bounded height on split toric varieties over $\mathbb{Q}$. Such varieties have proved to be a fertile testing ground for Manin’s conjecture and related conjectures, see e.g. [Sal98, Ess07, Hua22, PiSc23, PiSc24, Bon25].
@@
We give a precise asymptotic for this counting function as $B\rightarrow\infty$, establishing new cases of [Moe25conjecture, Conjecture 1.4].
@@
Here $a((X,M),L)$ and $b(\mathbb{Q},(X,M),L)$ are the Fujita invariant and the $b$-invariant of $(X,M)$ with respect to $L$ as defined in [Moe25conjecture], which we recall in Definition 2.17. Furthermore, if $L$ is adjoint rigid with respect to $(X,M)$, then the leading coefficient of $Q$ is the leading constant from [Moe25conjecture, Conjecture 1.4], and it is explicitly given in Theorem 3.3 as well as in Theorem 3.7.
Here the condition that $L$ is adjoint rigid means that the class $a((X,M),L)\operatorname{pr}_{M}^{*}L+K_{(X,M)}$ is represented by an unique $\mathbb{Q}$-divisor on $(X,M)$, see Definition 2.18. This theorem is a special case of Theorem 3.3, which only requires the pair to be quasi-proper as in Definition 2.23, rather than proper. To prove Theorem 1.1 for divisors which are not adjoint rigid with respect to $(X,M)$, we need to consider quasi-proper pairs to show that $Q$ has degree equal to, rather than at most, $b(\mathbb{Q},(X,M),L)-1$.
Theorem 1.1 implies [PSTVA21, Conjecture 1.1] for split toric varieties over $\mathbb{Q}$, including the conjecture for the leading constant if $L$ is adjoint rigid. This generalizes the results by Pieropan and Schindler [PiSc23, Theorem 1.2] to heights corresponding to divisors different from the log-anticanonical divisors, and improves on their error term. The theorem is proved using the universal torsor method, as developed by Salberger [Sal98].
The proof proceeds along the lines of de la Bretèche’s proof [Bre01Toric] of Manin’s conjecture for split toric varieties with the anticanonical height, together with Salberger’s computation of the leading constant [Sal98, Section 11]. Using the universal torsor, we identify the set of $\mathcal{M}$-points on $(\mathcal{X},\mathcal{M})$ with a subset of $\mathbb{Z}^{n}$, and we study the counting function via a related multivariate Dirichlet series. We then derive the asymptotic by applying two Tauberian theorems by de la Bretèche [Bre01Sum] to this series.
Our proof generalizes de la Bretèche’s proof [Bre01Toric] of Manin’s conjecture for split toric varieties in two ways: it generalizes generalizing his results to $\mathcal{M}$-points, and it furthermore works with other heights than the anticanonical height.
Theorem 1.1 is of interest even in the classical setting of rational points, as it improves on the original proof of Manin’s conjecture for toric varieties by Batyrev and Tschinkel [BaTs96, Corollary 1.5] by providing a good control of the error term. In the classical setting of rational points, Theorem 1.1 and its proof are similar to [Ess07, Theorem 1], where heights coming from other metrizations are considered. We finish the section by giving a few examples to illustrate Theorem 1.1. In these examples, we will consider weak Campana points.
@@
This result is related to Streeters result on powerful values of norm forms [Str21, Theorem 1.1], as the Campana pair considered in that article can be regarded as a Galois twist of $(\mathbb{P}^{n}_{\mathbb{Q}},\tfrac{1}{m}\sum_{i=1}^{n}D_{i})$. We discuss his result in further detail in [Moe25conjecture, §9.5].
More generally, Theorem 1.1 gives an asymptotic for the number of weak Campana points of bounded height on a split toric variety. See [Moe25conjecture, Theorem 1.7] for a direct description of the Fujita invariant and the $b$-invariant in this setting.
@@ 1.1. Structure of the paper
In Section 3, we recall the toric height used in this paper, and use it to state Theorem 3.3. This theorem gives an asymptotic formula for the counting function $N_{(X,M),L,S}$ for any quasi-proper pair $(X,M)$, and recovers Theorem 1.1 as a special case. Afterwards, we illustrate the theorem in a few examples and give a geometric interpretation of the leading constant.
Finally, we prove the Theorem 3.3 in two parts. In the first part we use [Bre01Sum, Théorème 1] to show the asymptotic $N_{(X,M),L,S}(B)=B^{a((X,M),L)}\left(Q(\log B)+O\left(B^{-\theta}\right)\right)$ for some polynomial $Q$ of degree at most $b(\mathbb{Q},(X,M),L)-1$ for which the holds. In the second part we use [Bre01Sum, Théorème 1] to show that $Q$ has degree exactly equal to $b(\mathbb{Q},(X,M),L)-1$, and we compute the leading coefficient under assumption that $L$ is toric adjoint rigid.
@@ 1.2. Acknowledgments
This article is based on my PhD thesis [Thesis], which I wrote at Utrecht University under supervision of Marta Pieropan. I would like to thank her for her unwavering support for my work and her various suggestions. I would also like to thank Alec Shute, Sam Streeter, Tim Santens and Damaris Schindler for the useful discussions we had during the workshop Campana points on toric varieties II in Bristol, as well as on other occasions. I would also like to thank Régis de la Bretèche for his thoughtful response to a question of mine. Finally, I would like to thank Sho Tanimoto for his useful comments on an earlier version of the paper.
@@ 1.3. Notation
For a field $K$ and a $\mathbb{Z}$-scheme $X$, we denote its base change to $K$ by $X_{K}=X\otimes_{\mathbb{Z}}\operatorname{Spec}K$.
@@ 2.1. Cox coordinates on split toric varieties
In this section we introduce smooth proper toric varieties and their Cox coordinates, which are a generalization of the homogeneous coordinates on projective space. We will use these Cox coordinates to describe and study $\mathcal{M}$-points on toric pairs. Let $N$ be a lattice and let $\Sigma$ be a regular and complete fan in $N$ as in [Sal98, Definition 8.1.1]. To such a fan, we define the associated toric scheme (over $\mathbb{Z}$) to be $\mathcal{X}_{\Sigma}$ as in [Sal98, Remark 8.6] and the associated toric variety over a field $K$ to be $X_{\Sigma}:=\mathcal{X}_{\Sigma,K}$. By construction, $\mathcal{X}_{\Sigma}$ is a proper and smooth $\mathbb{Z}$-scheme and $X:=X_{\Sigma}$ is a proper and smooth split toric variety over $K$.
@@
On a split toric variety, rational points can be described using Cox coordinates, generalizing the homogeneous coordinates on projective space. We recall the construction of Cox coordinates on such varieties and their integral model. For a regular and complete fan $\Sigma$ in $N\cong\mathbb{Z}^{d}$, [CLS11, Theorem 4.1.3] gives an exact sequence
@@
where the first map is the map sending a character to its corresponding divisor. When combined with [CLS11, Proposition 4.2.5], this sequence implies $\operatorname{Pic}(X)\cong\mathbb{Z}^{n-d}$. Furthermore, by this exact sequence, we can view the $\mathbb{Z}$-torus $\mathcal{G}:=\mathcal{H}\!\mathit{om}(\operatorname{Pic}(X),\mathbb{G}_{m,\mathbb{Z}})\cong\mathbb{G}_{m,\mathbb{Z}}^{n-d}$ as the kernel of the homomorphism
$\mathbb{G}_{m,\mathbb{Z}}^{n}\cong\mathcal{H}\!\mathit{om}(\mathbb{Z}^{\Sigma(1)},\mathbb{G}_{m,\mathbb{Z}})\rightarrow\mathbb{G}_{m,\mathbb{Z}}^{d}\cong\mathcal{H}\!\mathit{om}(N^{\vee},\mathbb{G}_{m,\mathbb{Z}})$
@@
$\mathcal{Z}=\{\mathbf{x}^{\hat{\sigma}}=0\mid\text{ for all }\sigma\in\Sigma_{\max}\}\subset\mathbb{A}_{\mathbb{Z}}^{n}.$
@@
of toric varieties. As in [CLS11, Theorem 5.1.11], this homomorphism is a $\mathcal{G}$-torsor, so $\pi$ gives an isomorphism
@@
In fact, $\pi$ is a universal torsor for $\mathcal{X}_{\Sigma}$. We use this isomorphism to define Cox coordinates. By [FrPi13, Proposition 2.1], any point $P\in\mathcal{X}_{\Sigma}(\mathbb{Z})$ is the image of a point $\tilde{P}\in\left(\mathbb{A}^{n}_{\mathbb{Z}}\setminus\mathcal{Z}\right)(\mathbb{Z}).$
@@
For a point $\tilde{P}\in\left(\mathbb{A}^{n}_{\mathbb{Z}}\setminus\mathcal{Z}\right)(\mathbb{Z})$, we say that $\tilde{P}=(a_{1},\dots,a_{n})$ are (integral) Cox coordinates for the point $P=\pi(\tilde{P})\in\mathcal{X}_{\Sigma}(\mathbb{Z})=X_{\Sigma}(\mathbb{Q})$. We will write $P=(a_{1}:\dots:a_{n})$, in analogy to the homogeneous coordinates on projective space. For $p$-adic points, we define Cox coordinates analogously.
@@ 2.2. Toric pairs and $\mathcal{M}$-points
A (smooth split) toric pair over a field $K$ is a divisorial pair $(X,M)$ as in [Moe25conjecture], where $X$ is a smooth proper split toric variety over $K$ and $M=((D_{1},\dots,D_{n}),\mathfrak{M})$, where $D_{1},\dots,D_{n}$ are the torus invariant prime divisors on $X$ and $\mathfrak{M}\subset\mathbb{N}^{n}$ is a subset containing the origin. The toric integral model of a toric pair $(X,M)$ is the the pair $(\mathcal{X},\mathcal{M})$ over $\mathbb{Z}$, where $\mathcal{X}$ is the toric integral model of $X$ and $\mathcal{M}=((\mathcal{D}_{1},\dots,\mathcal{D}_{n}),\mathfrak{M})$, where $\mathcal{D}_{1},\dots,\mathcal{D}_{n}$ are the Zariski closures of $D_{1},\dots,D_{n}$ in $\mathcal{X}$.
In contrast to [Moe25conjecture], we restrict $\mathfrak{M}$ to be a subset of $\mathbb{N}^{n}$ as we will only count rational points contained in the dense torus. Proper toric pairs are rationally connected by [Moe25conjecture, Proposition 5.3], so they satisfy the hypotheses of [Moe25conjecture, Conjecture 1.4].
@@
This multiplicity is the same as the multiplicity defined in [Moe25conjecture] determined by the divisors $\mathcal{D}_{1},\dots,\mathcal{D}_{n}$.

### v2

@@ Abstract.
We establish an asymptotic formula for the number of $\mathcal{M}$-points of bounded height on split toric varieties, for the height induced by any big and nef divisor class. This formula establishes new cases of the extension of Manin’s conjecture to $\mathcal{M}$-points, as introduced by the author. As a special case of our result, we strengthen the results obtained by Pieropan and Schindler on Campana points of bounded height on toric varieties. As another special case, we obtain an asymptotic for the number of weak Campana points of bounded height, which is novel even for projective space. We illustrate this result by giving an asymptotic for the number of points on projective space of bounded height for which the product of coordinates is powerful.
Finally, we derive an asymptotic for the number of rational points in the image of a toric rational map, in the spirit of the Loughran-Smeets conjecture.
@@ 1. Introduction
In this article, we obtain an asymptotic for the number of $\mathcal{M}$-points of bounded height on split toric varieties, in the spirit of Manin’s conjecture. Since its introduction in the 1980’s [16], Manin’s conjecture has inspired many developments in arithmetic geometry and arithmetic statistics, and it has received various refinements [4, 26, 7, 20]. This conjecture predicts a precise asymptotic for the number of rational points of bounded height on rationally connected varieties.
A few years ago, Manin’s conjecture was extended by Pieropan, Smeets, Tanimoto and Várilly-Alvarado to a conjecture on Campana points of bounded height [29]. This conjecture has attracted a lot of interest, and various cases of it have been proven [34, 35, 2, 27, 28, 11]. In a recent work [24], the author further extended Manin’s conjecture to the broader setting of $\mathcal{M}$-points. This conjecture gives a general framework for various results on Campana points and Darmon points which fall outside of the scope of the previous conjectures [32, 31, 19, 1]. In this article, we will prove the latter conjecture for $\mathcal{M}$-points of bounded height on split toric varieties over $\mathbb{Q}$. Such varieties have proved to be a fertile testing ground for Manin’s conjecture and related conjectures, see e.g. [30, 15, 18, 27, 28, 8].
@@
We give a precise asymptotic for this counting function as $B\rightarrow\infty$, establishing new cases of [24, Conjecture 1.4].
@@
Here $a((X,M),L)$ and $b(\mathbb{Q},(X,M),L)$ are the Fujita invariant and the $b$-invariant of $(X,M)$ with respect to $L$ as defined in [24], which we recall in Definition 2.21. Furthermore, if $L$ is adjoint rigid with respect to $(X,M)$, then the leading coefficient of $Q$ is the leading constant from [24, Conjecture 1.4], and it is explicitly given in Theorem 3.3 as well as in Theorem 3.8.
Here the condition that $L$ is adjoint rigid means that the class $a((X,M),L)\operatorname{pr}_{M}^{*}L+K_{(X,M)}$ is represented by a unique $\mathbb{Q}$-divisor on $(X,M)$, see Definition 2.22. This theorem is a special case of Theorem 3.3, which only requires the pair to be quasi-proper as in Definition 2.27, rather than proper. To prove Theorem 1.1 for divisors which are not adjoint rigid with respect to $(X,M)$, we need to consider quasi-proper pairs to show that $Q$ has degree equal to, rather than at most, $b(\mathbb{Q},(X,M),L)-1$.
Theorem 1.1 implies [29, Conjecture 1.1] for split toric varieties over $\mathbb{Q}$, including the conjecture for the leading constant if $L$ is adjoint rigid. This generalizes the results by Pieropan and Schindler [27, Theorem 1.2] to heights corresponding to divisors different from the log-anticanonical divisors, and improves on their error term. The theorem is proved using the universal torsor method, as developed by Salberger [30].
The proof proceeds along the lines of de la Bretèche’s proof [13] of Manin’s conjecture for split toric varieties with the anticanonical height, together with Salberger’s computation of the leading constant [30, Section 11]. Using the universal torsor, we identify the set of $\mathcal{M}$-points on $(\mathcal{X},\mathcal{M})$ with a subset of $\mathbb{Z}^{n}$, and we study the counting function via a related multivariate Dirichlet series. We then derive the asymptotic by applying two Tauberian theorems by de la Bretèche [14] to this series.
Our proof generalizes de la Bretèche’s proof [13] of Manin’s conjecture for split toric varieties in two ways: it generalizes generalizing his results to $\mathcal{M}$-points, and it furthermore works with other heights than the anticanonical height.
Theorem 1.1 is of interest even in the classical setting of rational points, as it improves on the original proof of Manin’s conjecture for toric varieties by Batyrev and Tschinkel [6, Corollary 1.5] by providing a good control of the error term. In the classical setting of rational points, Theorem 1.1 and its proof are similar to [15, Theorem 1], where heights coming from other metrizations are considered.
We now give a few examples to illustrate Theorem 1.1. In these examples, we will consider weak Campana points.
@@
This result is related to Streeters result on powerful values of norm forms [32, Theorem 1.1], as the Campana pair considered in that article can be regarded as a Galois twist of $(\mathbb{P}^{n}_{\mathbb{Q}},\tfrac{1}{m}\sum_{i=1}^{n}D_{i})$. We discuss his result in further detail in [24, §9.5].
More generally, Theorem 1.1 gives an asymptotic for the number of weak Campana points of bounded height on a split toric variety. See [24, Theorem 1.7] for a direct description of the Fujita invariant and the $b$-invariant in this setting.
As another consequence of Theorem 1.1, we also find an asymptotic formula for the number of rational points of bounded height in the image of a toric rational map, i.e. a rational map of toric varieties restricting to a homomorphism on the dense tori.
Corollary 1.5.
Let $f\colon Y\dashrightarrow X$ be a dominant toric rational map of smooth proper split toric varieties, let $(Y,M)$ be a proper toric pair with toric integral model $(\mathcal{Y},\mathcal{M})$, and let $L\in\operatorname{Pic}(X)$ be a big and nef divisor class. Then for all integers $B$, we have that
$\{P\in f((\mathcal{Y},\mathcal{M})(\mathbb{Z}))\mid H_{L}(P)\leq B\}=\Delta^{-1}N_{(X,f_{*}M),L,1}(B).$
Here $(X,f_{*}M)$ is the proper toric pair as in Definition 2.9, and $\Delta$ is the order of $(N/\overline{f}N^{\prime})\otimes\mathbb{Z}/2\mathbb{Z}$, where $\overline{f}\colon N^{\prime}\rightarrow N$ is the map of lattices corresponding to $f$.
This corollary is a direct consequence of Lemma 2.10 together with the fact that the height of a rational point $(a_{1},\dots:a_{n})$ does not depend on the signs of $a_{1},\dots,a_{n}$. Theorem 1.1 and Corollary 1.5 together give an explicit asymptotic for the number of rational points on $X$ of bounded height which lie in the image of an $\mathcal{M}$-point. In particular, this gives a formula for the number of rational points of bounded height in the image of $f$. This is in the spirit of the Loughran–Smeets conjecture and its extensions, as studied for instance in [21, 9]. It does not fall within the framework of the conjecture however, as the generic fiber of a toric morphism is usually not geometrically irreducible.
Example 1.6.
Let $f\colon\mathbb{P}^{2}\dashrightarrow\mathbb{P}^{2}$ be the rational map given by $(x:y:z)\mapsto(x^{2}:y^{2}:xz)$. Then the number of points in $f(\mathbb{P}^{2}(\mathbb{Q}))\subset\mathbb{P}^{2}(\mathbb{Q})$ of naive height at most $B$ tends to $B^{2}(Q(\log B)+O(B^{-\theta}))$, where $\theta>0$ and $Q$ is a linear polynomial with leading coefficient
$\frac{1}{2}\prod_{p\text{ prime}}(1-2p^{-2}+2p^{-3}-p^{-4})\approx 0.245.$
We will derive this asymptotic in Example 3.4.
@@ 1.1. Structure of the paper
In Section 3, we recall the toric height used in this paper, and use it to state Theorem 3.3. This theorem gives an asymptotic formula for the counting function $N_{(X,M),L,S}$ for any quasi-proper toric pair $(X,M)$, and recovers Theorem 1.1 as a special case. Afterwards, we illustrate the theorem in a few examples and give a geometric interpretation of the leading constant.
Finally, we prove the Theorem 3.3 in two parts. In the first part we use [14, Théorème 1] to show the asymptotic $N_{(X,M),L,S}(B)=B^{a((X,M),L)}\left(Q(\log B)+O\left(B^{-\theta}\right)\right)$ for some polynomial $Q$ of degree at most $b(\mathbb{Q},(X,M),L)-1$ for which the holds. In the second part we use [14, Théorème 1] to show that $Q$ has degree exactly equal to $b(\mathbb{Q},(X,M),L)-1$, and we compute the leading coefficient under assumption that $L$ is toric adjoint rigid.
@@ 1.2. Acknowledgments
This article is based on my PhD thesis [23], which I wrote at Utrecht University under the supervision of Marta Pieropan. I would like to thank her for her unwavering support for my work and her various suggestions. I would also like to thank Alec Shute, Sam Streeter, Tim Santens and Damaris Schindler for the useful discussions we had during the workshop Campana points on toric varieties II in Bristol, as well as on other occasions. I would also like to thank Régis de la Bretèche for his thoughtful response to a question of mine. Finally, I would like to thank Sho Tanimoto and Sam Streeter for their useful comments on an earlier version of the paper.
@@ 1.3. Notation
For a field $K$ and a $\mathbb{Z}$-scheme $X$, we denote its base change to $K$ by $X_{K}=X\times_{\mathbb{Z}}\operatorname{Spec}K$.
@@ 2.1. Cox coordinates on split toric varieties
In this section we introduce smooth proper toric varieties and their Cox coordinates, which are a generalization of the homogeneous coordinates on projective space. We will use these Cox coordinates to describe and study $\mathcal{M}$-points on toric pairs. Let $N$ be a lattice and let $\Sigma$ be a regular and complete fan in $N$ as in [30, Definition 8.1.1]. To such a fan, we define the associated toric scheme (over $\mathbb{Z}$) to be $\mathcal{X}_{\Sigma}$ as in [30, Remark 8.6] and the associated toric variety over a field $K$ to be $X_{\Sigma}:=\mathcal{X}_{\Sigma,K}$. By construction, $\mathcal{X}_{\Sigma}$ is a proper and smooth $\mathbb{Z}$-scheme and $X:=X_{\Sigma}$ is a proper and smooth split toric variety over $K$.
@@
On a split toric variety, rational points can be described using Cox coordinates, generalizing the homogeneous coordinates on projective space. We recall the construction of Cox coordinates on such varieties and their integral model. For a regular and complete fan $\Sigma$ in $N\cong\mathbb{Z}^{d}$, [12, Theorem 4.1.3] gives an exact sequence
@@
where the first map is the map sending a character to its corresponding divisor. When combined with [12, Proposition 4.2.5], this sequence implies $\operatorname{Pic}(X)\cong\mathbb{Z}^{n-d}$. Furthermore, by this exact sequence, we can view the $\mathbb{Z}$-torus $\mathcal{G}:=\operatorname{Hom}(\operatorname{Pic}(X),\mathbb{G}_{m,\mathbb{Z}})\cong\mathbb{G}_{m,\mathbb{Z}}^{n-d}$ as the kernel of the homomorphism
$\mathbb{G}_{m,\mathbb{Z}}^{n}\cong\operatorname{Hom}(\mathbb{Z}^{\Sigma(1)},\mathbb{G}_{m,\mathbb{Z}})\rightarrow\mathbb{G}_{m,\mathbb{Z}}^{d}\cong\mathcal{H}\!\mathit{om}(N^{\vee},\mathbb{G}_{m,\mathbb{Z}})$
@@
$\mathcal{Z}=\{(x_{1},\dots,x_{n})\in\mathbb{A}_{\mathbb{Z}}\mid\mathbf{x}^{\hat{\sigma}}=0\text{ for all }\sigma\in\Sigma_{\max}\}.$
@@
of toric varieties. As in [12, Theorem 5.1.11], this homomorphism is a $\mathcal{G}$-torsor, so $\pi$ gives an isomorphism
@@
In fact, $\pi$ is a universal torsor for $\mathcal{X}_{\Sigma}$. We use this isomorphism to define Cox coordinates. By [17, Proposition 2.1], any point $P\in\mathcal{X}_{\Sigma}(\mathbb{Z})$ is the image of a point $\tilde{P}\in\left(\mathbb{A}^{n}_{\mathbb{Z}}\setminus\mathcal{Z}\right)(\mathbb{Z}).$
@@
For a point $\tilde{P}=(a_{1},\dots,a_{n})\in\left(\mathbb{A}^{n}_{\mathbb{Z}}\setminus\mathcal{Z}\right)(\mathbb{Z})$, we say that $(a_{1},\dots,a_{n})$ are (integral) Cox coordinates for the point $P=\pi(\tilde{P})\in\mathcal{X}_{\Sigma}(\mathbb{Z})=X_{\Sigma}(\mathbb{Q})$. We will write $P=(a_{1}:\dots:a_{n})$, in analogy to the homogeneous coordinates on projective space. For $p$-adic points, we define Cox coordinates analogously.
@@ 2.2. Toric pairs and $\mathcal{M}$-points
A (smooth split) toric pair over a field $K$ is a divisorial pair $(X,M)$ as in [24], where $X$ is a smooth proper split toric variety over $K$ and $M=((D_{1},\dots,D_{n}),\mathfrak{M})$, where $D_{1},\dots,D_{n}$ are the torus invariant prime divisors on $X$ and $\mathfrak{M}\subset\mathbb{N}^{n}$ is a subset containing the origin. The toric integral model of a toric pair $(X,M)$ is the the pair $(\mathcal{X},\mathcal{M})$ over $\mathbb{Z}$, where $\mathcal{X}$ is the toric integral model of $X$ and $\mathcal{M}=((\mathcal{D}_{1},\dots,\mathcal{D}_{n}),\mathfrak{M})$, where $\mathcal{D}_{1},\dots,\mathcal{D}_{n}$ are the Zariski closures of $D_{1},\dots,D_{n}$ in $\mathcal{X}$.
In contrast to [24], we restrict $\mathfrak{M}$ to be a subset of $\mathbb{N}^{n}$ as we will only count rational points contained in the dense torus. Proper toric pairs are rationally connected by [24, Proposition 5.3], so they satisfy the hypotheses of [24, Conjecture 1.4].
@@
This multiplicity is the same as the multiplicity defined in [24, Definition 3.8] determined by the divisors $\mathcal{D}_{1},\dots,\mathcal{D}_{n}$.
@@
2.3. Images of $\mathcal{M}$-points under toric rational maps
We now describe how images of rational points under toric rational maps can be viewed as $\mathcal{M}$-points on a toric pair. For this we first recall the homomorphism $\phi$ from [22, Definition 5.10].
Definition 2.8.
Let $X$ be a smooth split toric variety with torus invariant prime divisors $D_{1},\dots,D_{n}$ and corresponding ray generators $n_{\rho_{1}},\dots,n_{\rho_{n}}$. We define $\phi\colon\mathbb{Z}^{n}\rightarrow N$ to be the group homomorphism given by
$(m_{1},\dots,m_{n})\rightarrow\sum_{i=1}^{n}m_{i}n_{\rho_{i}}.$
Let $f\colon X\dashrightarrow X^{\prime}$ be a toric rational map of smooth proper split toric varieties over $\mathbb{Q}$, i.e. a homomorphism $U\rightarrow U^{\prime}$ of the dense tori in $X$ and $X^{\prime}$, and let $\overline{f}\colon N\rightarrow N^{\prime}$ be the corresponding map of lattices.
Definition 2.9.
For a toric pair $(X,M)$, let $(X^{\prime},f_{*}M)$ be the toric pair given by the set of multiplicities
$f_{*}\mathfrak{M}=(\phi^{\prime})^{-1}(\overline{f}\circ\phi)(\mathfrak{M}_{\textnormal{red}}),$
where $\phi\colon\mathbb{Z}^{n}\rightarrow N,\phi^{\prime}\colon\mathbb{Z}^{n^{\prime}}\rightarrow N^{\prime}$ are the homomorphisms from Definition 2.8 for $X$ and $X^{\prime}$.
The images of $\mathcal{M}$-points under $f$ are $f_{*}\mathcal{M}$-points, and the converse holds up to changing the signs of the coordinates.
Lemma 2.10.
Let $f\colon X\dashrightarrow X^{\prime}$ be a toric rational map of smooth split toric varieties over $\mathbb{Q}$ and let $(X,M)$ be a toric pair. Let $\mathcal{U},\mathcal{U}^{\prime}$ be the dense tori of $\mathcal{X}$ and $\mathcal{X}^{\prime}$. Then
$f((\mathcal{X},\mathcal{M})(\mathbb{Z}))\subset(\mathcal{X}^{\prime},f_{*}\mathcal{M})(\mathbb{Z}).$
Furthermore, for every $P\in(\mathcal{X}^{\prime},f_{*}\mathcal{M})(\mathbb{Z})$, there exists $u\in\mathcal{U}^{\prime}(\mathbb{Z})$ such that the product $uP$ lies in the image, and $u$ is unique modulo $f(\mathcal{U}(\mathbb{Z}))$.
Proof.
Let $p$ be a prime number. The composition $\phi\circ\operatorname{mult}_{p}\colon U(\mathbb{Q}_{p})\rightarrow N$ is a surjective group homomorphism with kernel $\mathcal{U}(\mathbb{Z}_{p})$, see e.g. [22, Proposition 5.12.], giving a natural isomorphism $N\cong U(\mathbb{Q}_{p})/\mathcal{U}(\mathbb{Z}_{p})$. Therefore, the diagram
[…]
commutes. Therefore for every $Q\in(\mathcal{X},\mathcal{M})(\mathbb{Z}_{p})$, we have that $\operatorname{mult}_{p}(f(Q))\in f_{*}\mathfrak{M}$. Furthermore, the Chinese remainder theorem implies that for every $P\in(\mathcal{X}^{\prime},f_{*}\mathcal{M})(\mathbb{Z})$ there is $Q\in(\mathcal{X},\mathcal{M})(\mathbb{Z})$ such that $f(Q)$ has the same multiplicities as $P$ at all prime numbers. Therefore, they differ by an $\mathbb{Z}$-integral point $u$ on the torus $\mathcal{U}^{\prime}$. Finally, if is another integral point $u^{\prime}$ such that $u^{\prime}P$ lies in the image, then $u^{-1}u^{\prime}$ also lies in the image. ∎
Remark 2.11.
Lemma 2.10 generalizes from $\mathbb{Z}$ to any ring of $\mathcal{O}_{K,S}$ integers, over a global field $K$, provided $S$ is chosen large enough to ensure that $\mathcal{O}_{K,S}$ is a principal ideal domain. This is done by replacing the role of the Chinese remainder theorem in the proof with the fact that $\mathbb{A}^{1}$ satisfies strong approximation away from the infinite places.
Example 2.12.
Let $X$ be a smooth proper split toric variety over $\mathbb{Q}$, let $m$ be a positive integer, and let $f\colon X\rightarrow X$ be the morphism corresponding to the multiplication-by-$m$ map $N\rightarrow N$. Then $f(X(\mathbb{Q}))$ are the set of points in $X(\mathbb{Q})$ of the form $(a_{1}^{m}:\dots:a_{n}^{m})$, for nonzero rational numbers $a_{1},\dots,a_{n}$. On the other hand, if $(X,M)$ is the pair given by $\mathfrak{M}=\mathbb{N}^{n}$, then $(\mathcal{X},f_{*}\mathcal{M})(\mathbb{Z})$ are the strict Darmon points corresponding to the Campana pair $\left(X,\sum_{i=1}^{n}\left(1-\tfrac{1}{m}\right)D_{i}\right)$ (as defined in [22, Definition 3.16]). These are the points of the form $(\pm a_{1}^{m}:\dots:\pm a_{n}^{m})$, for nonzero rational numbers $a_{1},\dots,a_{n}$. In particular we see that the set of Darmon points is equal to the image exactly when $m$ is odd.